# Incorporating Geo-Diverse Knowledge into Prompting for Increased Geographical Robustness in Object Recognition

### v1

@@ Abstract
Existing object recognition models have been shown to lack robustness in diverse geographical scenarios due to significant domain shifts in design and context. Class representations need to be adapted to more accurately reflect an object concept under these shifts. In the absence of training data from target geographies, we hypothesize that geography-specific descriptive knowledge of object categories can be leveraged to enhance robustness. For this purpose, we explore the feasibility of probing a large-language model for geography-specific object knowledge, and we investigate integrating knowledge in zero-shot and learnable soft prompting with the CLIP vision-language model. In particular, we propose a geography knowledge regularization method to ensure that soft prompts trained on a source set of geographies generalize to an unseen target set of geographies. Our gains on DollarStreet when generalizing from a model trained only on data from Europe are as large as +2.8 on countries from Africa, and +4.6 on the hardest classes. We further show competitive performance vs. few-shot target training, and provide insights into how descriptive knowledge captures geographical differences.
@@ 1 Introduction
The performance of object recognition models has been shown to degrade when tested on geographies (e.g., cities, countries, continents) that differ from training data [7, 39, 21, 33, 43]. Numerous factors contribute to the challenging problem of geographical domain shift, such as cross-geography changes in object design/parts, materials, and context. These changes in turn may be due to cultural, climate, and economic differences across the world. Recent work has shown standard adaptation techniques fail when used for geographical domain shifts [21, 33], but there has yet to be significant progress in the creation of techniques that improve geographical robustness. Such progress is necessary to ensure equitable use of AI as use grows.
@@
Figure 1: Descriptive knowledge can bridge concept shifts across geographies. Observe the wide range of object designs and contexts in the DollarStreet [11] category tools around the world. Our work’s premise is that textual representations for classes in vision-language models can be tailored to better suit diverse object representations across geographies. Map made with [16].
@@
Overall, models need representations that adequately capture a category’s various forms around the world. A natural solution is collecting training data of objects from different regions. However, this approach is expensive, takes significant effort, and is difficult for low-resource geographies. Fortunately, geographical shifts have a unique property compared to other common domain shifts (e.g. ones due to artistic style or weather changes)—they can be addressed with descriptive knowledge about concept changes. In other words, it is possible to describe the more representative features of an object in a geography and to use this information to adapt a model’s default representation. For instance, as shown in Fig. 1, areas in Papua New Guinea may have tools often tailored towards “cooking, hunting, and fishing”, and areas in Malawi may have tools often made out of “stone or wood” material. Models should account for geo-diverse interpretations of a category and not be limited to biased interpretations (e.g. if the model learns tools as just being “metallic with logos”).
@@
We explore the feasibility of leveraging geo-diverse knowledge in two ways. First, we examine whether a vision-language model (VLM, i.e. CLIP [36]) has encoded categories in a geography-specific manner, such that adding a country’s name to a prompt (e.g. “A photo of tools in Malawi”) elicits knowledge that improves recognition. Second, we probe a large language model (LLM, i.e. GPT-3 davinci-003) for geography-specific knowledge to obtain representative visual feature descriptors for an object across the world. We identify complementary strengths and combine VLM- and LLM-based knowledge for zero-shot inference on socioeconomically diverse data (DollarStreet [11]).
@@
We further consider a practical scenario where we optimize CLIP with soft prompting, using only a “source” geography with easy-to-access data (e.g. Europe), while the models are applied downstream on “target” data from other parts of the world (e.g. Africa, Asia, Americas). We propose geography knowledge regularization, which uses knowledge ensembled over countries to enable soft prompts to achieve geographically generalizable class representations. We test our methods on the recent DollarStreet and GeoNet [21] datasets. Our regularization boosts performance over baseline soft prompting methods, and has benefits with respect to few-shot target-specific training (a 16-shot-per-class regularized model without any target data outperforms a 12-shot-per-class target-trained model on DollarStreet).
@@
Our method is the first that effectively addresses geographical shifts in object recognition. It outperforms zero-shot CLIP (which may be assumed to have some robustness) by 10.3% on Africa, CoOp by 3.3%, and the best baseline by 4.6% on the hardest classes.
@@
To summarize, we answer the following questions: (1) Does adding geographical context (i.e., country names) to CLIP prompts improve recognition across geographies? (2) Can an LLM provide useful geographical descriptive knowledge to improve recognition? (3) How can we optimize soft prompts for CLIP using an accessible data source with consideration of target geographies not represented in the training set? (4) Where can soft prompts enhanced with geographical knowledge provide the most benefits?
@@
Figure 2: Geography knowledge regularization. To ensure robustness in soft prompt learning, we (1) incorporate complementary knowledge internal to CLIP and externally obtained from an LLM. (2) This descriptive knowledge regularizes class representations when training on a specific source geography (e.g. Europe), thus (3) increasing robustness when generalizing to unseen geographies (e.g. Togo).
@@ 2 Related Work
Geographical domain shifts occur when the target setting is in a different geography (e.g. continent, country, city) than where the source data was acquired. Shifts involve changes in object design (e.g. differences in house architecture) and context (i.e. background/co-occurring objects vary). Datasets tailored to cross-country/continent object recognition have recently been proposed, e.g. DollarStreet [11], GeoNet [21], GeoDE [37], GeoYFCC [9], and OpenImages-Extended [5]. Interestingly, [33, 21] demonstrate that traditional methods in unsupervised domain adaptation [10, 27, 28, 19, 45, 50, 38, 44, 20] which seek to bridge gaps based on visual features alone, do not effectively address geographical domain shift. They achieve negligible gains (e.g. 0.14 for [10] in [33]) or often drops in performance (all methods tested in [21]), compared to just using the source model. Attempts to specifically address geographical robustness are limited: [43] corrects for differences in the sizes of cars, [9] proposes a discriminative domain embedding from target data, and GiVL [48] pretrains with knowledge from Wikipedia. In contrast, our descriptive knowledge regularization works for different categories (not just cars); we do not require target domain data to achieve gains cross-geography; we leverage the strong capabilities of LLMs to gather relevant knowledge; and we propose lightweight adaptation through zero-shot and soft prompting (unlike GiVL’s expensive pretraining).
@@
Vision-language (VL) models [36, 17, 25, 26, 49] excel on a variety of tasks. CLIP [36] shows impressive zero-shot object recognition across different settings. Yet its performance given geographical shift is less apparent, with GeoNet only showing finetuned performance, which is expensive given CLIP’s large scale. [37] on the other hand only show zero-shot with default CLIP prompts. Neither method considers descriptive knowledge or soft prompting.
@@
Learning soft textual prompts. Several recent works to adapt CLIP have focused on parameter and data efficiency using linear probing [36] and prompting [52, 18, 23]. Soft textual prompting (e.g. CoOp [52]) is notable as it optimizes class text embeddings (without manual tuning), which we hypothesize is critical to adequately adapt for geographical robustness. As CoOp overfits on base (seen) classes, CoCoOp [51] proposes to condition prompts on the image for better generalizability. KgCoOp [47] alternatively guides learned prompt embeddings towards CLIP’s manual prompt embeddings through a distance constraint to avoid degradation on unseen classes. Our approach also uses a distance constraint, but it differs from [47] with the purpose of regularizing learned prompt representations for cross-geography generalization instead of the base-to-new-class setting. We also show novel benefits of regularization when used with an ensemble of CLIP’s internal geographical knowledge and external geographical descriptive knowledge. Our approach notably outperforms each of CoOp, CoCoOp, and KgCoOp by at least +2.8 accuracy on unseen countries in Africa in DollarStreet. External knowledge aids unseen classes in KAPT [22], but not with respect to geographical knowledge. Prompt tuning for adaptation has been tested in [40, 12], but not with descriptive knowledge.
@@
Probing knowledge from large language models like [4, 31, 32, 6] has been used for visual reasoning [46], embodied agent planning [15, 41], and to generate additional context for class prompts to VLMs in object recognition [30, 34]. We uniquely probe LLMs for distinguishable visual descriptions for the same object class across different geographical regions. We are also the first to incorporate geographical knowledge from LLMs into soft prompting.
@@ 3 Approach
We address geographical shift in object recognition with VLMs. We posit that the manner in which classes are described is critical due to cross-geography shifts in design and context. We hypothesize that CLIP’s default class representations elicited through “a photo of a/an $<$object$>$” prompts may not adequately represent classes around the world. Instead, they may be more aligned to high-resource geographies due to Internet-based training data. Optimizing representations (with soft prompts) on a specific geography (e.g. Europe) may exacerbate a lack of robustness. Our main idea (Fig. 2) is to incorporate object-related geographical knowledge into prompting to ensure model robustness with respect to diverse regions and cultures. We outline a recipe to obtain geography-specific context by probing CLIP’s internal knowledge and an external LLM’s descriptive knowledge. We further propose geography knowledge regularization to ensure soft prompts do not overfit when training data is limited to high-resource geographies.
Preliminaries. We consider object recognition on a dataset $\mathcal{S}$ containing a class set $\mathcal{C}$ (size $N_{c}$) over across a set of geographies $\mathcal{G}$. We consider a geography $g$ to be either a country or continent. Our VLM is CLIP [36], with an image encoder $f$ and language encoder $h$. We incorporate knowledge of each geography $g$ into prompting using: (1) zero-shot inference or (2) soft textual prompting. Prompts are defined as $\boldsymbol{t}$ (set of tokens), and class embeddings $\boldsymbol{w}$ are calculated as $h(\boldsymbol{t})$. We refer to CLIP’s default prompt “a photo of a/an $<$object$>$” for a class $c$ as $\boldsymbol{t}_{c}^{\text{default}}$.
@@ 3.1 Geographical Knowledge Probing
Probing CLIP’s internal geographical knowledge. Our first strategy is to augment CLIP’s manual prompts to include country names, as we surmise the resulting class representations may be better aligned to how objects present in diverse geographies. [3] inspires this hypothesis, showing that adding country names to image generation prompts achieves significant gains in geographical representativeness. However, it is an open question whether adding country names in prompts improves recognition. We define the setting CountryInPrompt, using the prompt $\boldsymbol{t}_{c}^{\text{CountryInPrompt}}$ with template “a photo of a/an $<$object$>$ in $<$country$>$”, e.g. “a photo of a stove in Burundi.”
Leveraging external LLM geographical knowledge. As CLIP may not have sufficient knowledge of objects in various countries, we consider further augmenting prompts with external knowledge. Motivated by probing LLMs for general attribute-based object descriptions [30, 34] (e.g. a tiger with “stripes and sharp teeth”), we probe GPT-3 (davinci-003) for geography-specific descriptions of object styles, contexts, and materials. [^1] We found ChatGPT to perform worse than GPT-3, also found in [34]. We reason that since LLMs are typically trained on large information sources (e.g. CommonCrawl [1], WebText [35], Wikipedia [2]) they may have knowledge about how an object presents in a region due to climate, economics, and other cultural factors. For instance, roofs may be “thatched” in tropical and temperate climates, and cutlery may made of “bamboo” in areas with bamboo forests. Our goal is unique vs. [30, 34] in that we aim to use descriptive knowledge differences for the same class to address domain shifts across regions.
Acquiring knowledge. We follow [30], but instead of gathering one set of feature descriptors $\mathcal{D}(c)$ for each $c$, we collect a set per country. For each class $c$ and geography $g$, we prompt the LLM with following prompt template, consisting of an example question, answer, and format, to generate descriptor lists $\mathcal{D}_{g}(c)$. We use 1-shot prompting (give an example in the prompt) to capture geographically representative object designs and context, as exemplified with the descriptors for Japanese ofuro (お風呂, bathtub) below:
@@
Using knowledge. To convert LLM outputs to CLIP prompts, each descriptor $d$ in $\mathcal{D}_{g}(c)$ serves as a prompt $\boldsymbol{t}$. The format is of $\boldsymbol{t}$ is “a photo of a/an $<$object$>$ which (is/has/etc.) $<$descriptor$>$”. The setting where geography-specific LLM descriptors are used in CLIP prompting is referred to as CountryLLM (prompts $\boldsymbol{t}_{c,d}^{\text{CountryLLM}}$), while [30] is GeneralLLM (prompts $\boldsymbol{t}_{c,d}^{\text{GeneralLLM}}$). To perform zero-shot inference on an image $I$, each class score $s(c,I)$ is computed using the average CLIP logits, $\phi(I,d)$, over each $d$ in the set $\mathcal{D}$. For GeneralLLM, the score is calculated as:

### v2

@@ Abstract
Existing object recognition models have been shown to lack robustness in diverse geographical scenarios due to domain shifts in design and context. Class representations need to be adapted to more accurately reflect an object concept under these shifts. In the absence of training data from target geographies, we hypothesize that geographically diverse descriptive knowledge of categories can enhance robustness. For this purpose, we explore the feasibility of probing a large language model for geography-based object knowledge, and we examine the effects of integrating knowledge into zero-shot and learnable soft prompting with CLIP. Within this exploration, we propose geography knowledge regularization to ensure that soft prompts trained on a source set of geographies generalize to an unseen target set. Accuracy gains over prompting baselines on DollarStreet while training only on Europe data are up to +2.8/1.2/1.6 on target data from Africa/Asia/Americas, and +4.6 overall on the hardest classes. Competitive performance is shown vs. few-shot target training, and analysis is provided to direct future study of geographical robustness.
@@ 1 Introduction
The performance of object recognition models degrades when tested in new geographies (e.g., cities, countries, continents) [7, 39, 21, 33, 43]. Numerous factors contribute to the challenging problem of geographical domain shift, such as cross-geography changes in object design/parts, materials, and context. These changes in turn may be due to cultural, climate, or economic differences around the world. Recent work has shown standard adaptation techniques fail when used for geographical domain shifts [21, 33], but there has yet to be significant progress in the creation of techniques that improve geographical robustness. Such progress is necessary to ensure equitable use of AI in the future.
@@
Figure 1: Descriptive knowledge can address concept shifts across geographies. Observe the wide range of object designs and contexts in the DollarStreet [11] category tools around the world. Our work’s premise is that textual representations for classes in vision-language models can be enhanced to better suit diverse object representations across geographies. Map made with [16].
@@
Overall, models need representations that adequately capture a category’s various forms around the world. A natural solution is to collect training data of objects from different regions. However, this approach is expensive, takes significant effort, and is difficult for regions with limited Internet access. Fortunately, geographical shifts have a unique property compared to other common domain shifts (e.g. ones due to artistic style or weather changes)—they can be addressed with descriptive knowledge about concept changes. In other words, it is possible to describe the features of an object in a region and use this information to adapt a model’s default representation. For instance, as shown in Fig. 1, for rural areas in Papua New Guinea, tools can be described as being used for “cooking, hunting, and fishing”, and for rural areas in Malawi, tools may often be “made of metal and wood, for farming”. Models should account for diverse presentations and contexts of a category and not be limited to biased presentations (e.g. if the model learns tools as just being “metallic with logos”).
@@
We examine the effects of probing geo-diverse knowledge in two ways. First, we analyze whether a vision-language model (VLM, i.e. CLIP [36]) has encoded categories in a geo-specific manner, such that adding a country’s name to a prompt (e.g. “A photo of a house in China”) elicits knowledge that improves recognition. Second, we probe a large language model (LLM, i.e. GPT-3 davinci-003) for geography-specific knowledge to obtain visual feature descriptors for an object in different locations. We analyze results in zero-shot inference on geographically and socioeconomically diverse data (DollarStreet [11]), finding the combination of knowledge to often be complementary.
@@
We further consider a practical scenario where CLIP is optimized with soft prompting, using only a “source” geography with easy-to-access data (e.g. Europe), while the model is applied downstream on “target” data from other parts of the world (e.g. Africa, Asia, Americas). We propose geography knowledge regularization, which uses knowledge ensembled over countries to enable soft prompts to achieve geographically generalizable class representations. We test our method on the recent DollarStreet and GeoNet [21] datasets. Our regularization boosts performance over baseline soft prompting methods, and has benefits with respect to few-shot target-specific training (a 16-shot-per-class regularized model without any target data outperforms a 12-shot-per-class target-trained model on DollarStreet).
@@
Our method is the first to effectively address geo shifts in object recognition. It outperforms zero-shot CLIP (assumed to have some robustness) by 10.3% on Africa, CoOp [52] by 3.3%, and the best baseline by 4.6% on the hardest classes.
@@
To summarize, we answer the following questions: (1) Does adding geographical context (i.e. country names) to CLIP prompts improve recognition across geographies? (2) Can an LLM provide useful geographical descriptive knowledge to improve recognition? (3) How can we optimize soft prompts for CLIP using an accessible data source with consideration of target geographies not represented in the training set? (4) Where can soft prompts enhanced with geographical knowledge provide the most benefits?
@@
Figure 2: Geography knowledge regularization. To ensure robustness in soft prompt learning, we (1) incorporate knowledge internal to CLIP and externally obtained from an LLM. (2) This descriptive knowledge regularizes class representations when training on a specific source geography (e.g. Europe), thus (3) increasing robustness when generalizing to target geographies (e.g. Vietnam).
@@ 2 Related Work
Geographical domain shifts occur when the target setting is in a different geography (e.g. continent, country, city) than where the source data was acquired. Shifts involve changes in object design (e.g. differences in house architecture) and context (i.e. background/co-occurring objects vary). Datasets tailored to cross-country/continent object recognition have recently been proposed, e.g. DollarStreet [11], GeoNet [21], GeoDE [37], GeoYFCC [9], and OpenImages-Extended [5]. Interestingly, [33, 21] demonstrate that traditional methods in unsupervised domain adaptation [10, 27, 28, 19, 45, 50, 38, 44, 20] which seek to bridge gaps based on visual features alone, do not effectively address geographical domain shift. They achieve negligible gains (e.g. 0.14 for [10] in [33]) or often drops in performance (e.g. all methods tested in [21]), compared to just using the source model. Attempts to specifically address geographical robustness are limited: [43] corrects for differences in the sizes of cars, [9] proposes a discriminative domain embedding from target data, and GiVL [48] pretrains with knowledge from Wikipedia. In contrast, our descriptive knowledge regularization works for different categories (not just cars); we do not require target domain data to achieve gains cross-geography; we explore the strong capabilities of LLMs to gather relevant knowledge; and we propose lightweight adaptation through soft prompting (unlike GiVL’s expensive pretraining).
@@
Vision-language (VL) models [36, 17, 25, 26, 49] excel on a variety of tasks. CLIP [36] shows impressive zero-shot object recognition across different settings. Yet its performance given geographical shift is less apparent. GeoNet [21] only shows finetuned performance, which is expensive given CLIP’s large scale. GeoDE [37] only shows zero-shot inference with CLIP’s default prompts. Neither work evaluates descriptive knowledge or soft prompting.
@@
Learning soft textual prompts. Several recent works to adapt CLIP have focused on parameter and data efficiency using linear probing [36] and prompting [52, 18, 23]. Soft textual prompting (e.g. CoOp [52]) is notable as it optimizes class text embeddings (without manual tuning), which we hypothesize is critical to adequately adapt for geographical robustness. As CoOp overfits on base (seen) classes, CoCoOp [51] proposes to condition prompts on the image for better generalizability. KgCoOp [47] alternatively guides learned prompt embeddings towards CLIP’s manual prompt embeddings through a distance constraint to avoid degradation on unseen classes. Our approach also uses a distance constraint, but it differs from [47] with the purpose of regularizing learned prompt representations for cross-geography generalization instead of the base-to-new-class setting. We also show novel benefits of regularization when used with an ensemble of CLIP’s internal geographical knowledge and external geographical descriptive knowledge. Our approach notably outperforms each of CoOp, CoCoOp, and KgCoOp by at least +2.8 accuracy on target countries in Africa in DollarStreet. External knowledge aids unseen classes in KAPT [22], but not with respect to geographical knowledge. Prompt tuning for adaptation has been tested in [40, 12], but not with descriptive knowledge.
@@
Knowledge probed from large language models like [4, 31, 32, 6] has been used for visual reasoning [46], embodied agent planning [15, 41], and to generate additional context for VLM class prompts in object recognition [30, 34]. We uniquely probe LLMs for distinguishable visual descriptions for the same object class across different geographical regions. We are also the first to incorporate geographical knowledge from LLMs into soft prompting.
@@ 3 Approach
We investigate geographical shift in object recognition with VLMs. We posit that the manner in which classes are described is critical due to cross-geography shifts in design and context. We also hypothesize that CLIP’s default class representations elicited through “a photo of a/an $<$object$>$” prompts may not adequately represent classes around the world. Instead, they may be more aligned to high-resource geographies due to Internet-based training data. Optimizing representations (with soft prompts) on a specific geography (e.g. Europe) may exacerbate a lack of robustness. Our main idea (Fig. 2) is to incorporate object-related geographical knowledge into prompting to ensure model robustness in different regions. We outline our mechanism to obtain geography-specific context by probing CLIP’s internal knowledge and an external LLM’s descriptive knowledge. We further propose geography knowledge regularization to ensure soft prompts do not overfit when training data is limited to certain geographies.
Preliminaries. We consider object recognition on a dataset $\mathcal{S}$ containing a class set $\mathcal{C}$ (size $N_{c}$) over a set of geographies $\mathcal{G}$. We consider a geography $g$ to be either a country or continent. Our VLM is CLIP [36], with an image encoder $f$ and language encoder $h$. We incorporate knowledge of each geography $g$ into prompting using (1) zero-shot inference or (2) soft textual prompting. Prompts are defined as $\boldsymbol{t}$ (each is a set of tokens), and class embeddings $\boldsymbol{w}$ are calculated as $h(\boldsymbol{t})$. We refer to CLIP’s default prompt “a photo of a/an $<$object$>$” for a class $c$ as $\boldsymbol{t}_{c}^{\text{default}}$.
@@ 3.1 Geographical Knowledge Probing
Probing CLIP’s internal geographical knowledge. Our first strategy of investigation is to augment CLIP’s manual prompts to include country names, as we surmise that some of the resulting class representations may be better aligned to how categories present in different regions. [3] inspires this hypothesis, showing that adding country names to image generation prompts can achieve gains in geographical representativeness. However, it is an open question whether adding country names in prompts improves recognition. We define the setting CountryInPrompt, using the prompt $\boldsymbol{t}_{c}^{\text{CountryInPrompt}}$ with template “a photo of a/an $<$object$>$ in $<$country$>$”, e.g. “a photo of a stove in Burundi.”
Probing external LLM geographical knowledge. As CLIP may not have sufficient knowledge of objects in some regions, we consider further augmenting prompts with external knowledge. Motivated by probing LLMs for general attribute-based object descriptions [30, 34] (e.g. a tiger with “stripes and sharp teeth”), we probe GPT-3 (davinci-003) for geography-specific descriptions of object styles, contexts, and materials. [^1] We found ChatGPT to perform worse than GPT-3, also found in [34]. We reason that since LLMs are trained on large information sources (e.g. CommonCrawl [1], WebText [35], Wikipedia [2]), they may have knowledge about how an object presents in a region due to climate, economics, and/or cultural factors. For instance, roofs may sometimes be “thatched” in tropical and temperate climates, and cutlery may sometimes be made of “bamboo” in areas with bamboo forests. Our goal is unique vs. [30, 34] in that we explore descriptive knowledge differences for the same class to address domain shifts across regions.
Acquiring knowledge. We follow [30], but instead of gathering one set of feature descriptors $\mathcal{D}(c)$ for each $c$, we collect sets per country. For each class $c$ and geography $g$, we prompt the LLM to generate descriptor lists $\mathcal{D}_{g}(c)$, using a template consisting of an example question, answer, and format. We use 1-shot prompting to show how to capture geographically representative object designs and contexts. Our prompt exemplifies this below, using the descriptors for Japanese ofuro{CJK}UTF8min (お風呂, bathtub):
@@
Using knowledge. To convert LLM outputs to CLIP prompts, each descriptor $d$ in $\mathcal{D}_{g}(c)$ serves in a prompt $\boldsymbol{t}_{c,d}$. The format of $\boldsymbol{t}_{c,d}$ is “a photo of a/an $<$object$>$ which (is/has/etc.) $<$descriptor$>$". The setting where geography-specific LLM descriptors are used in prompting is referred to as CountryLLM (prompts $\boldsymbol{t}_{c,d}^{\text{CountryLLM}}$), while [30] is GeneralLLM (prompts $\boldsymbol{t}_{c,d}^{\text{GeneralLLM}}$). To perform zero-shot inference on an image $I$, each class score $s(c,I)$ is computed using the average of CLIP logits $\phi(I,d)$ over each $d$ in the set $\mathcal{D}$. For GeneralLLM, the score is calculated as:
$s(c,I)=\frac{1}{|\mathcal{D}(\textit{c})|}\sum_{d\in\mathcal{D}(\textit{c})}\phi(I,d)$ (1)
For CountryLLM, we use the geo-specific set:
$s(c,I,g)=\frac{1}{|\mathcal{D}_{g}(\textit{c})|}\sum_{d\in\mathcal{D}_{g}(\textit{c})}\phi(I,d)$ (2)
The argmax of $s$ with respect to $c$ is taken as the prediction. Due to averaging over descriptor scores, not every descriptor needs to strongly activate in a correct prediction. The model therefore can account for diverse features of objects within a geography. These descriptors effectively serve as complements to CLIP’s default knowledge of class names.
Combining knowledge. Our third method of exploration, CountryInPrompt+LLM, combines both CLIP’s internal knowledge and LLM external knowledge. The prompt template $\boldsymbol{t}_{c,d}^{\text{CountryInPrompt+LLM}}$ is “a photo of a/an $<$object$>$ in $<$country$>$ which (is/has/etc.) $<$descriptor$>$".
3.2 Regularizing Soft Prompts via Geo Knowledge
Adaptation scenario. In practice, one may want to further optimize a VLM for a downstream task. To update a model effectively, one promising strategy is soft textual prompting. It is parameter-efficient [52] and avoids feature distortion unlike finetuning [24]. Its mechanism is to learn context parameters that directly change the class text embeddings used in inference. We posit that learning context on a dataset with limited diversity (e.g. just Europe) may tailor these class representations to the region and overfit. To investigate cross-geography generalization when using soft prompting, we pose a domain generalization scenario where we aim to learn only from a high-resource source set of countries and generalize to a target set of countries at inference time. A method that performs well in this setting could provide a viable alternative to few-shot target training when acquiring target data for training is not feasible.
Soft prompts. Our idea is to learn soft prompts while constraining the class text embeddings to be close to geographical knowledge of objects outside of source geographies. In this way, we hope to learn class representations that are more applicable to the rest of the world. Building from CoOp [52], we assume there is a text prompt $\boldsymbol{t}_{c}$ for each class $c$. All prompts share $M$ context vectors (each denoted [V]m), which are the same size as the word embeddings (i.e. 512-D) and precede a class name token [CLASSc]: